### v1

# Zero-Shot Image Segmentation via Recursive Normalized Cut on Diffusion Features

Abstract
Foundation models have emerged as powerful tools across various domains including language, vision, and multimodal tasks. While prior works have addressed unsupervised image segmentation, they significantly lag behind supervised models. In this paper, we use a diffusion UNet encoder as a foundation vision encoder and introduce DiffCut, an unsupervised zero-shot segmentation method that solely harnesses the output features from the final self-attention block. Through extensive experimentation, we demonstrate that the utilization of these diffusion features in a graph based segmentation algorithm, significantly outperforms previous state-of-the-art methods on zero-shot segmentation.  Specifically, we leverage a recursive Normalized Cut algorithm that softly regulates the granularity of detected objects and produces well-defined segmentation maps that precisely capture intricate image details. Our work highlights the remarkably accurate semantic knowledge embedded within diffusion UNet encoders that could then serve as foundation vision encoders for downstream tasks. Project page: https://diffcut-segmentation.github.io
Figure 1: Unsupervised zero-shot image segmentation. Our DiffCut method  exploits features from a diffusion UNet encoder in a graph-based recursive partitioning algorithm. Compared to DiffSeg [1],  DiffCut provides finely detailed segmentation maps that more closely align with semantic concepts.
1 Introduction
Foundation models have emerged as powerful tools across various domains, including language [2, 3, 4], vision [5, 6, 7], and multimodal tasks [8, 9, 10, 11, 12, 13]. Pretrained on extensive datasets, these models exhibit unparalleled generalization capabilities, marking a significant departure from training models from scratch to efficiently adapting pretrained foundation models [14, 15, 16, 17]. Utilizing pretrained models is particularly vital for dense visual tasks, alleviating the need for large annotated datasets specific to each domain.  While prior works [18, 19, 20, 21, 22] have addressed unsupervised image segmentation, they significantly lag behind supervised models [23, 24, 25, 26].  Recently, SAM [27], proposed a model that can produce fine-grained class-agnostic masks which achieves outstanding zero-shot transfer to any images. Still, it requires a huge annotated segmentation dataset as well as significant training resources. Therefore, in this work, we investigate an alternative direction: unsupervised and zero-shot segmentation under the most constraining  conditions, where no segmentation annotations or prior knowledge on the target dataset are available.
Recently, several methods have emerged to address unsupervised object detection by framing it as a graph partitioning problem, utilizing self-supervised ViT features [28, 5]. LOST [29] proposes to localize a unique object in a image by exploiting the inverse degree information to find a seed patch. TokenCut [30] splits the graph in two subsets given a bipartition. FOUND [31] and MaskCut [32] extend these approaches by addressing the single object discovery limitation. While being able to localize multiple objects, the latter methods remain constrained to identify a pre-determined number of objects, making them ill-suited for a task of unsupervised image segmentation which inherently requires to adapt the number of segment to uncover to the visual content.
Conversely, text-to-image diffusion models [33, 34, 35] can produce high-quality visual content from textual descriptions [36, 37, 38], indicating implicit learning of a wide range of visual concepts.  Recent works have tried to leverage diverse internal representations of such models for localization or segmentation tasks. Several methods [39, 40, 41, 42, 43] opt to exploit image-text interactions within cross-attention modules but are ultimately constrained by the need for meticulous input prompt design. Concurrently, [44] identifies semantic correspondences between image pixels and spatial locations of low-dimensional feature maps by modulating cross-attention modules. This method proves to be computationally intensive as it requires numerous forward inferences. On the other hand, DiffSeg [1] segment images by iteratively merging self-attention maps which only depict local correlation between patches.
In this work, we introduce DiffCut, a new method for zero-shot image segmentation which solely harnesses the encoder features of a pre-trained diffusion model in a recursive graph partitioning algorithm  to produce fine-grained segmentation maps. Importantly, our method does not require any label from downstream segmentation datasets and its backbone has not been pre-trained on dense pixel annotations such as SAM [27].  We observe in Fig. 1 that DiffCut produces sharp segments that nicely outline object boundaries.  In comparison with the recent state-of-the-art unsupervised zero-shot segmentation method DiffSeg [1], the segments yielded by DiffCut, are better aligned with the semantic visual concepts, e.g. DiffCut is able to uncover the urban area as well as the boats in the middle row image. Our main contributions are as follows:
  • We leverage the features from the final self-attention block of a diffusion UNet encoder, for the task of unsupervised image segmentation. In this context, we demonstrate that exploiting the inner patch-level alignment yields superior performance compared to merging self-attention maps as done in DiffSeg [1].
  • Compared to existing graph based object localization methods e.g. TokenCut or MaskCut [30, 32], we push further and take advantage of a recursive Normalized Cut algorithm to generate dense segmentation maps. Via a partitioning threshold, the method is able to regulate the granularity of detected objects and consequently adapt the number of segments to the visual content.
  • We perform extensive experiments to validate the effectiveness of DiffCut and show that it significantly outperforms state-of-the-art methods for unsupervised segmentation on standard benchmarks, reducing the gap with fully supervised models.
In addition, we exhibit the remarkable semantic coherence emerging in our chosen diffusion features by measuring their patch-level alignment, which surpasses other backbones such as CLIP [8] or DINOv2 [5]. Our ablation studies further reveal the relevance of these diffusion features as well as the recursive partitioning approach which proves to provide robust segmentation performance. Finally, we show that DiffCut can be extended to an open-vocabulary setting with a straightforward process leveraging a convolutional CLIP, which even tops most dedicated methods on this task.
2 Related Work
Semantic segmentation.
Semantic segmentation consists in partitioning an image into a set of segments, each corresponding to a specific semantic concept. While supervised semantic segmentation  has been widely explored [45, 46, 47, 27], unsupervised and zero-shot transfer segmentation for any images with previously unseen categories remains significantly more challenging and much less investigated.  For example, most works in unsupervised segmentation require access to the target data for unsupervised adaption [21, 20, 19, 18]. Therefore, these methods cannot segment images that are not seen during the adaptation. Recently, DiffSeg [1] moved a step forward by proposing an unsupervised and zero-shot approach that can produce quality segmentation maps without any prior knowledge on the underlying visual content.
Segmentation with Text Supervision.
Recent works have shown that learning accurate segmentation maps is possible with text supervision, overcoming the cost of dense annotations. These works are mostly based on image-text contrastive learning [48, 49, 50, 51], and usually exploit the features of CLIP [52, 53, 54]. MaskCLIP [52] leverages CLIP to get pseudo labels used to train a typical image segmentation model. ReCO [53] uses CLIP for dataset curation and get a reference image embedding for each class that is used to obtain the final segmentation. CLIPpy [48] proposes minimal modifications to CLIP to get dense labels. SegCLIP [54] continues to train CLIP with additional reconstruction and superpixel-based KL loss to enhance localization. TCL [50] learns a region-text alignment to get precise segmentation masks. GroupViT [49] also learns masks from text supervision and is based on a hierarchical grouping mechanism. Similarly, ViewCo [51] proposes a contrastive learning between multiple views/crops of the image and the text.
Graph-based Object Detection.
Built on top of self-supervised ViT features, various methods frame the problem of object detection as a graph partitioning problem. LOST [29] aims at detecting salient object in an image using the degree of the nodes in the graph and a seed expansion mechanism. Based on Normalized Cut (NCut) [55], FOUND [31] proposes to identify all background patches, hence discovering all object patches as a by-product with no need for a prior knowledge of the number of objects or their relative size with respect to the background. TokenCut [30] detects one single salient object in each image with a unique NCut bipartition. In an attempt to adapt TokenCut to multi-objects localization, MaskCut [32] first localizes an object and disconnects its corresponding patches to the rest of the graph before repeating the process a pre-determined number of times.  As these graph partitioning methods are only able to uncover a fixed number of segments, they are inadequate for a task of image segmentation.
Segmentation with Diffusion Models.
Diffusion models can produce high-quality visual content given a text prompt, indicating implicit learning of a wide range of visual concepts and the ability of grounding these concepts in images. Therefore their internal representations appear as good candidates for visual localization tasks [56, 57, 58].  ODISE [59] is one of the first training-based approaches to build a fully supervised panoptic image segmentor on top of diffusion features. Several other methods [40, 41, 42] leverage attention modules for localization or segmentation tasks. DiffuMask [42] uses the cross-modal grounding between a text input and an image in cross-attention modules to segment the referred object in a synthetic image. However, DiffuMask can only be applied to a generated image.  In a zero-shot setting, [41] harnesses the image-text interaction via cross-attention score maps to complete self-attention maps and segment grounded objects. EmerDiff [44] opts not to exploit image-text interactions in cross-attention modules. Instead, it identifies semantic correspondences between image pixels and spatial locations by modulating the values of a sub-region of feature maps in low-resolution cross-attention layers. These cross-attention based methods eventually prove to be highly computationally intensive as multiple forward inferences are often required. On the other hand, DiffSeg [1] proposes an iterative merging process based on measuring KL divergence among self-attention maps to merge them into valid segmentation masks. However, it appears that self-attention score maps only depict very local correlation between patches.
3 DiffCut
Diffusion Models.
Diffusion models [60, 61, 62] are generative models that aim to approximate a data distribution $q$ by mapping an input noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$ to a clean sample $\bm{x}_{0}\sim q$ through an iterative denoising process. In latent text-to-image (T2I) diffusion models, e.g. Stable Diffusion [33], the diffusion process is performed in the latent space of a Variational AutoEncoder [63] for computational efficiency, and encode the textual inputs as feature vectors from pretrained language models. Starting from a noised latent vector $\mathbf{z_{t}}$ at the timestep $t$, a denoising autoencoder $\epsilon_{\theta}$ is trained to predict the noise $\epsilon$ that is added to the latent $\mathbf{z}$, conditioned on the text prompt $\mathbf{c}$. The training objective writes:
$\mathcal{L}=\mathbb{E}_{\mathbf{z}\sim\mathcal{E}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,1),t}\left[\lVert\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\tau(\mathbf{c}))\rVert^{2}_{2}\right]$ (1)
where $t$ is uniformly sampled from the set of timesteps $\{1,\ldots,T\}$.
Figure 2: Overview of DiffCut. 1) DiffCut takes an image as input and extracts the features of the last self-attention block of a diffusion UNet encoder. 2) These features are used to construct an affinity matrix that serves in a recursive normalized cut algorithm, which outputs a segmentation map at the latent spatial resolution. 3) A high-resolution segmentation map is produced via a concept assignment mechanism on the features upsampled at the original image size.
3.1 Features Extraction
An input image  is encoded into a latent via the VQ-encoder of the latent diffusion model and a small amount of gaussian noise is added to it (not shown in Fig. 2). The obtained latent is passed to the diffusion UNet encoder, from which we only extract the output features, denoted $\hat{z}$, from its last self-attention block.  This choice design has several motivations:
Attention Limitations.
In contrast to several methods that harness cross-attention modules for localization or segmentation tasks [39, 40, 41, 42], we deliberately choose not to depend on this mechanism. The accuracy of segmentation maps generated via attention modules heavily relies on the quality of the textual input which often requires an automatic captioning model combined with a meticulous prompt design to reach competitive performance. Besides being constrained by the maximum number of input tokens, such approach is proved to be inaccurate in the presence of cohyponyms [64] and is prone to neglect subject tokens as the number of objects to detect becomes large [65]. The localization and segmentation capacity with a single forward inference is then constrained by the performance of the captioning model and the attention modules themselves. Exploiting only the intermediate diffusion features alleviate the computational cost of an additional captioning model and do not necessitate multiple forward inferences.
UNet Encoder Effectiveness.
Previous works [66, 67, 37] have shown that diffusion features provide precise semantic information shared across objects from different domains. Building on this observation, we hypothesize that the pyramidal architecture of the UNet encoder capture semantically rich image representations that are well-suited for zero-shot vision tasks. To validate this assumption, we exhibit  the semantic coherence emerging in the UNet encoder, evidenced by a remarkable patch level alignment in the output features of the final self-attention block. We in fact demonstrate  that these features are sufficient to reach state-of-the-art zero-shot segmentation performance.
Computational Efficiency.
By solely exploiting the diffusion UNet encoder, our method offers a substantial computational gain, reducing the model size by $70\%$ (400M vs 1.3B parameters). In contrast, DiffSeg extracts every self-attention maps of the UNet which requires a full model inference.
3.2 Recursive Feature Clustering
Normalized Cut treats image segmentation as a graph partitioning problem [55]. Given a graph $\mathbf{G=(V,E)}$ where $\mathbf{V}$ and $\mathbf{E}$ are respectively a set of nodes and edges, we construct an affinity matrix $\bm{W}$ such that $\bm{W}_{ij}$ is the edge between node $v_{i}$ and $v_{j}$, and a diagonal degree matrix $\bm{D}$, with $d(i)=\sum_{j}\bm{W}_{ij}$. NCut minimizes the cost of partitioning the graph into two sub-graphs by solving:
$(\bm{D-W})x=\lambda\bm{D}x$ (2)
to find the eigenvector $x$ corresponding to the second smallest eigenvalue. In the ideal case, the clustering solution only takes two discrete values.  Since the solution of Eq. 2 is a continuous relaxation of the initial problem, $x$ contains continuous values and a splitting point has to be determined to partition it.  To find the optimal partition, we examine $l$ evenly spaced points in $x$ and select the one resulting in the minimum NCut value.
Graph Affinity.
Deriving from our observation on the patch-level alignment of our chosen diffusion features, we assume that the normalized cut algorithm would provide sharp segments, each corresponding to a precise semantic concept as distinct objects would manifest as weakly connected components in a patch similarity matrix. Following this intuition, we construct an affinity matrix $\bm{W}$, by computing the cosine similarity between each pair of patches.  In addition, since NCut criterion assesses both the overall dissimilarity among different segments and the similarity within each segment, we opt to emphasize inter-segments dissimilarity by exponentiating each element with a positive integer value $\alpha$:
$\bm{W}_{ij}=\left(\frac{\hat{z}_{i}\hat{z}_{j}}{\lVert\hat{z}_{i}\rVert_{2}\lVert\hat{z}_{j}\rVert_{2}}\right)^{\alpha}$ (3)
Essentially, this process maintains a relatively high affinity for highly similar patches, while squashing the weights between dissimilar patches towards zero. This mechanism plays the role of a soft thresholding, offering a more gradual adjustment compared to setting a threshold to explicitly binarize the affinity matrix as done in [30] and [32].
Recursive Partitioning.
Classical spectral clustering [68] requires setting a pre-defined number of clusters to partition the graph, which is a significant constraint in the context of zero-shot image segmentation where no prior knowledge on the visual content is available. We therefore adopt a recursive graph partitioning [55], which adapts the number of uncovered segments to the visual content via a threshold, denoted $\tau$, on the maximum partitioning cost.  This hyperparameter stops the recursive partitioning of a segment when its NCut value exceeds it and thereby regulates the granularity of detected objects. We demonstrate that our soft thresholding process detailed above enhances the robustness of the method which delivers competitive performance across a wide range of $\tau$ values. This recursive clustering process is summarized in Supplementary A.
3.3 High-Resolution Concept Assignment
Thus far, we have constructed segmentation maps (e.g. $32\times 32$) which are $32$ times lower in resolution than the original image (e.g. $1024\times 1024$). The number of segments found in each image depends on the image and the value of the hyperparameter $\tau$. Our next goal is to upscale these low-resolution maps to build accurate pixel-level segmentation maps. We propose a high-resolution segmentation process that can be decomposed into the following steps:
  1. 1.
    Masked Spatial Marginal Mean. First, our objective is to extract a set of representations that embeds the semantics of each segment. As shown in [37], reducing the spatial dimension of diffusion features with a Spatial Marginal Mean (SMM) effectively retains semantic information and provides accurate image descriptor. In light of this, we naturally propose to collapse the spatial dimension of each segment with a Masked SMM. This process yields a collection of semantically rich concept-embeddings, denoted $\mathcal{C}$.
  2. 2.
    Concept Assignment. A naive approach to obtain segmentation maps at the original image resolution consists in performing a nearest-neighbor upsampling. Despite its straightforwardness, this approach results in a blocky output structure as all pixels within the same feature patch are assigned to the same concept. Alternatively, we opt to first bilinearly upsample our low-resolution feature map $\hat{z}$ to match the original image spatial size and then proceed with the pixel/concept assignment. Specifically, for each concept $c_{i}\in\mathcal{C}$, we compute its cosine similarity with the upsampled features $\hat{z}_{up}$. This yields a similarity matrix of size $(H\times W\times K)$ where $K=|\mathcal{C}|$. Then, the assignment process simply consists in taking the argmax across the $K$ channels of the similarity matrix. The obtained segmentation map $\mathcal{S}$ is eventually refined with a pixel-adaptive refinement module [69].
4 Experiments
Datasets.
Following existing works in image segmentation [21, 53, 52, 18], we use the following datasets for evaluation: a) Pascal VOC [70] (20 foreground classes), b) Pascal Context [71] (59 foreground classes), c) COCO-Object [72] (80 foreground classes), d) COCO-Stuff-27 merges the 80 things and 91 stuff categories in COCO-stuff into 27 mid-level categories, e) Cityscapes [73] (27 foreground classes) and f) ADE20K (150 foreground classes) [74]. An extra background class is considered in Pascal VOC, Pascal Context, and COCO-Object. We ignore their training sets and directly evaluate our method on the original validation sets, at the exception of COCO for which we evaluate on the validation split curated by prior works [21, 19].
Metrics.
For all datasets, we report the mean intersection over union (mIoU), the most popular evaluation metric for semantic segmentation. Because our method does not provide a semantic label, we use the Hungarian matching algorithm [75] to assign predicted masks to a ground truth mask. For datasets including a background class, we perform a many-to-one matching to the background label. As in [1], we emphasize unsupervised adaptation (UA), language dependency (LD), and auxiliary image (AX). UA means that the specific method requires unsupervised training on the target dataset. This is common in the unsupervised segmentation literature. Methods without the UA requirement are considered zero-shot. LD means that the method requires text input, such as a descriptive sentence for the image, to facilitate segmentation. AX means that the method requires an additional pool of reference images or synthetic images.
Implementation details.
DiffCut builds on SSD-1B [35], a distilled version of Stable Diffusion XL [34]. The model takes an empty string as input and we set the timestep for denoising to $t=50$. To ensure a fair comparison when evaluating our method against baselines, we set a unique value for $\tau$ and $\alpha$ across all datasets, equal to $0.5$ and $10$ respectively.  Following previous works, we make use of PAMR [69] to refine our segmentation masks. Our method runs on a single NVIDIA TITAN RTX (24GB) with input images of size $1024\times 1024$ and can segment an image in one second.
4.1 Results on Zero-shot Segmentation
Tab. 1 reports the mIoU score for each baseline across the 6 benchmarks. Note that the numbers shown for COCO-Stuff and Cityscapes are taken from [1].  We complete ReCo [53] and MaskCLIP [52] scores with the results obtained in [50].  Other numbers are taken from [18].  We also note that DiffSeg tunes the sensible merging hyperparameter on a subset of images from the training set from the respective datasets. For a fair comparison, we evaluate the method fixing it to 1, as recommended in the original paper, and refine the obtained masks with PAMR. This baseline is denoted DiffSeg.
Table 1: Unsupervised segmentation results. Best method in bold, second is underlined.
Model LD AX UA VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K Extra-Training    IIC [19] 9.8 - - 6.7 6.4 -    MDC [76] - - - 9.8 7.1 -    PiCIE [21] - - - 13.8 12.3 -    PiCIE+H [21] - - - 14.4 -    STEGO [20] - - - 28.2 21.0 -    ACSeg [18] 53.9 - - 28.1 - - Training-free    ReCO [53] 25.1 19.9 15.7 26.3 19.3 11.2    MaskCLIP [52] 38.8 23.6 20.6 19.6 10.0 9.8    DiffSeg [1] - - - 43.6 21.2 -    DiffSeg 49.8 48.8 23.2 44.2 16.8 37.7    DiffCut (Ours) 65.2 56.5 34.1 49.1 30.6 44.3
With our set of default hyperparameters, DiffCut significantly outperforms all other baselines despite not relying on language dependency, auxiliary images or unsupervised adaptation. In average, our method gains +8.5 mIoU over the second best baseline. In particular, it significantly surpasses the previous state-of-the-art method in unsupervised segmentation, DiffSeg,  by +5.5 and +9.4 mIoU on COCO-Stuff and Cityscapes respectively. The effectiveness of our method is further corroborated with our qualitative results shown in Fig. 1. In comparison to DiffSeg, DiffCut provides finely detailed segmentation maps that more closely align with semantic concepts. Additional examples can be found in Supplementary J.
We note here that, as the granularity of annotations varies across target datasets, our fixed set of hyperparameters can not be in the optimal regime on each of them. Therefore, relaxing the condition on prior knowledge about the target dataset, we report in Supplementary E results of DiffCut where $\tau$ is loosely tuned using a small subset of annotated images from the target training split.
4.2 Semantic Coherence in Vision Encoders
As good candidates for a task of unsupervised segmentation are expected to be semantically coherent, we conduct a comparison between different families of foundation models on their internal alignment at the patch-level. Selected models include text-to-image DMs (SSD-1B [35]), text-aligned contrastive models (CLIP [77], SigLIP [78]) and self-supervised models (DINO [28], DINOv2 [5]). At the exception of DINO-ViT-B/16, evaluated models are of roughly similar size, approximately 300M parameters for DINOv2, CLIP-ViT-L/14 and SigLIP-ViT-L/16 and 400M for SSD-1B UNet encoder.
Figure 3: ROC curves revealing the semantic coherence of vision encoders.
As in [79], we collect patch representations from various vision encoders and store their corresponding target classes using the segmentation labels. Given $\hat{z}_{i}=\mathcal{E}(\mathbf{x_{1}})_{i}\in\mathbb{R}^{D_{v}}$ and $\hat{z}_{j}=\mathcal{E}(\mathbf{x_{2}})_{j}\in\mathbb{R}^{D_{v}}$, the patch representations of images $\mathbf{x_{1}}$ and $\mathbf{x_{2}}$ at respectively index $i$ and $j$, we compute their cosine similarity  and use this score as a binary classifier to predict if the two patches belong to the same class. Given $l(\mathbf{x_{1}})_{i}$ and $l(\mathbf{x_{2}})_{j}$, the labels associated to the patches, if $l(\mathbf{x_{1}})_{i,j}=l(\mathbf{x_{2}})_{p,q}$, the target value for binary classification is 1, else 0. We present in Fig. 3 the ROC curve and AUC score for our candidate models. We observe that SSD-1B UNet encoder [35] demonstrates a greater patch-level alignment than any other candidate model with an AUC score of 0.83, even surpassing DINOv2 [5]. We further exhibit the outstanding alignment between patch representations associated to semantically similar concepts with qualitative results in Fig. 4. We provide additional qualitative examples patch-level alignment in Supplementary I.
Ref. Image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption SSD-1B (ours) DINO-VIT-B/16 DINOv2-VIT-L/14 CLIP-VIT-L/14 SigLIP-VIT-L/16
Figure 4: Qualitative results on the semantic coherence of various vision encoders. We select a patch (red marker) associated to the dog in Ref. image. Top row shows the cosine similarity heatmap between the selected patch and all patches produced by vision encoders for Ref. image. Bottom row shows the heatmap between the selected patch in Ref. image and all patches produced by vision encoders for Target. image.
A potential rationale for this observation lies in the superior semantic information retention of a diffusion model compared to alternative backbones, attributed to its inherent capacity to set a structural image layout, internally acquired during the training phase. These results provides insight into the strong clustering results presented in previous section, as improved semantic coherence suggests that patches belonging to the same object are more effectively clustered.
4.3 Ablation study
In this section, we perform ablation studies to validate the individual choices in the design of DiffCut.
DiffCut vs DiffSeg.
DiffSeg proposes their own clustering algorithm based on a self-attention map merging process. As the original implementation uses a different diffusion backbone as ours,  we validate the benefit of our method by swapping the original SDv1.4 with our stronger SSD-1B. For a fair comparison between methods, we use the default set of hyperparameters recommended in [1] and set the default merging threshold of DiffSeg to $0.5$ for all datasets. Tab. 2 clearly validates the superiority of using rich semantic features in a recursive graph partitioning algorithm over the self-attention merging mechanism of DiffSeg. Qualitative results shown in Fig. 1 further display the edge of DiffCut in uncovering semantic clusters. Shown results do not make use of the mask refinement module, explaining the gap with Tab. 1.
Table 2: Ablation Study. DiffCut recursive partitioning algorithm yields superior results than self-attention merging of DiffSeg and Automated Spectral Clustering.
Model VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K DiffSeg 48.2 41.2 31.7 35.4 22.3 39.9 AutoSC 61.5 53.3 29.8 46.9 25.3 38.9 DiffCut (w/o PAMR) 62.0 54.1 32.0 46.1 28.4 42.4
DiffCut vs Automated Spectral Clustering.
In DiffCut, the hyperparameter $\tau$ corresponds to the maximum graph partitioning cost allowed.  In contrast, classical spectral clustering requires to explicitly set the number of segments to be found in the graph. To validate the benefit of the recursive approach over spectral clustering, we introduce a simple yet effective baseline dubbed AutoSC. [80] proposes a heuristic that estimates the number of connected components in a graph with the largest relative-eigen-gap between its Laplacian eigen-values. The larger the gap, the more confident the heuristic. In our context, the index of the eigen-value that maximizes this quantity can be interpreted as the number of clusters in an image. Thus, we define a set of exponents $\{1,5,10,15\}$ and determine the value $\alpha$ in this set such that its element-wise exponentiation of matrix $\bm{A}$ yields the largest Laplacian relative-eigen-gap. Then, we use the index of the eigen-value maximizing the gap as the number of clusters in a $k$-way spectral clustering performed with the algorithm proposed in [81]. As shown in Tab. 2, DiffCut consistently outperforms AutoSC on all datasets, with a gain up to $+3.5$ on ADE20K, at the exception of COCO-Stuff where the latter yields slightly better results. Noting that AutoSC is already a state-of-the-art baseline  on most benchmarks, this experiment confirms the relevance of the recursive Normalized Cut to uncover arbitrary numbers of segments.
4.4 Model Analysis
Figure 5: Robustness of DiffCut. As $\alpha$ increases, DiffCut shows competitive results for a broad range of $\tau$ values.
Segmentation Robustness.
In this section, we assess the impact of hyperparameters $\tau$ and $\alpha$ over the segmentation performance. We report in Fig. 5 the mIoU for various $\alpha$ values, with respect to partitioning threshold values $\tau$ ranging from 0.3 to 0.97 on Cityscapes validation set. As $\alpha$ increases, we observe a dual effect. First, since a greater $\alpha$ value shrinks the affinity matrix components towards 0, the partitioning cost corresponding to the NCut value decreases, explaining the shift of the optimal threshold between the different curves. Second, as $\alpha$ increases, the range of $\tau$ values for which the method yields competitive performance widens, contributing to the overall robustness of the method. For example, DiffCut outperforms our own competitive baseline AutoSC for any $\tau$ between 0.35 and 0.67 when $\alpha=10$, whereas it only surpasses it between 0.92 and 0.96 when $\alpha=1$.  Qualitatively, we observe in Fig. 6 that as $\tau$ increases, the method uncovers finer segments in images.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Ref. Image $\tau=0.3$ $\tau=0.5$ $\tau=0.7$
Figure 6: Effect of $\tau$. As $\tau$ corresponds to the maximum Ncut value, a larger threshold loosens the constraint on the partitioning algorithm and allows it to perform more recursive steps to uncover finer objects. It can be interpreted as the level of granularity of detected objects.
Table 3: Features Contribution. Hierarchical features in $\mathcal{E}_{32}$ provide optimal performance (Pascal VOC validation set).
$\mathcal{E}$ $\mathcal{D}$ VOC Test 32 64 32 64 mIoU - - - 62.0 - - 61.6 60.9
Diffusion Features.
Our chosen diffusion backbone uses a UNet-based architecture which consists of an encoder $\mathcal{E}$, bottleneck $\mathcal{B}$ and decoder $\mathcal{D}$. The hierarchical features of the encoder, with spatial resolution of $128\times 128$, $64\times 64$ and $32\times 32$ respectively, are injected into the decoder $\mathcal{D}$ via skip connections. Considering the final self-attention modules at resolution $64\times 64$ and $32\times 32$ in the encoder and decoder, we demonstrate in Tab. 3, that the encoder’s features extracted at the lowest spatial resolution retain the most semantic information and are sufficient to reach optimal performance. In addition, combining different hierarchical features does not lead to any improvements and adds up to the computational burden.
Table 4: Open-Vocabulary Segmentation. A straightforward open-vocabulary extension with a CNN-based CLIP yields competitive performance.
Model LD VOC Context COCO-Object Extra-Training    ReCO [53] 25.1 19.9 15.7    ViL-Seg [82] 37.3 18.9 -    MaskCLIP [52] 38.8 23.6 20.6    TCL [50] 55.0 30.4 31.6    CLIPpy [48] 52.2 - 32.0    GroupVIT [49] 52.3 22.4 24.3    ViewCo [51] 52.4 23.0 23.5    SegCLIP [54] 52.6 24.7 26.5    OVSegmentor [83] 53.8 20.4 25.1 Training-free    CLIP-DIY [84] 59.9 19.7 31.0    FreeSeg-Diff [85] 53.3 - 31.0    DiffCut 63.0 24.6 36.0
Open-Vocabulary Extension.
In the optic to extend DiffCut to an open-vocabulary setting, we propose in Tab. 4, a straightforward extension to associate a semantic label to each segmentation mask. Our proposed extension reaches competitive performance, even outperforming most baselines dedicated to the task of open-vocabulary zero-shot semantic segmentation. To do so, we leverage a frozen convolutional CLIP backbone and perform a mask-pooling over its semantic features to borrow its pretrained open-vocabulary recognition ability. Once mask proposals are predicted, they are classified with category text embeddings in a contrastive manner, where the class embeddings for each mask and category text embeddings are projected into a common embedding space.
5 Discussion
In this work, we address the challenging task of unsupervised and zero-shot image segmentation by introducing DiffCut, which significantly narrows the performance gap with fully supervised models. Specifically, DiffCut leverages diffusion features of a UNet encoder in a recursive graph partitioning algorithm to produce segmentation maps and reaches state-of-the-art performance on popular benchmarks. Yet, fully supervised end-to-end segmentation approaches still have an edge in terms of both performance and efficiency. Additional developments to further bridge the gap with supervised models could pave the way to use diffusion UNet encoders as foundation vision encoders for future downstream tasks.
Acknowledgement
We thank Louis Serrano, Adel Nabli, and Louis Fournier for their insightful discussions and helpful suggestions on the paper.
Supplementary Material
[…]
Appendix A Recursive Normalized Cut on Diffusion Features
We summarize in Algorithm 1 the recursive clustering process used in DiffCut.
Algorithm 1 Recursive Normalized Cut on Diffusion Features
Input:  $\mathcal{I}$ an image, $\tau\in]0,1[$ a threshold value, $\alpha\in\mathbb{N}^{*}$ an exponent value.
Step 1: Features Extraction.
  • Encode the image $\mathcal{I}$ with the VQ-encoder of the diffusion model: $z=\mathcal{E}_{\text{VAE}}(\mathcal{I})$
  • Add some gaussian noise to the latent image $z$.
  • Pass the noisy latent to the diffusion UNet encoder and extract the output features from its last self-attention block: $\hat{z}=\mathcal{E}_{\text{UNet}}(z)$
Step 2: Graph Construction.
  • Compute the pairwise cosine-similarity between patches of $\hat{z}$ to set up a similarity matrix $\bm{A}$.
  • Raise to the power $\alpha$ each element of matrix $\bm{A}$ to obtain the affinity matrix $\bm{W}$ (see Eq. 3).
  • Determine the matrix degree $\bm{D}$.
Step 3: NCut Problem Solving.
  • Solve $(\bm{D-W})x=\lambda\bm{D}x$ for eigenvector with the second smallest eigenvalue.
  • Use the eigenvector with the second smallest eigenvalue to bipartition the graph by finding the splitting point such that the NCut value is minimized.
Step 4: Recursive Partitioning.
  • Store the current partition and retrieve the matrices $\bm{W}$ and $\bm{D}$ associated to each sub-graph.
  • Recursively subdivide the partitions (Step 3) until the NCut value is greater than $\tau$.
Output:$M$ a segmentation map with the spatial resolution of $\hat{z}$.
Appendix B Impact of PAMR
To reveal the effect of the pixel-adaptive refinement module (PAMR) on our method, we compare the segmentation results on all benchmarks both with and without it enabled.
Table 5: Impact of PAMR on unsupervised segmentation.
DiffCut VOC Context COCO-Object COCO-Stuff Cityscapes ADE20K w/o PAMR 62.0 54.1 32.0 46.1 28.4 42.4 w/ PAMR 65.2 56.5 34.1 49.1 30.6 44.3
In average, PAMR allows to gain +2.5 mIoU on our segmentation benchmarks. Even though this refinement module helps to better outline the contour of objects, our method still reaches state-of-the-art results on unsupervised zero-shot segmentation without it.
Appendix C Mask Upsampling
Table 6: Mask Upsampling.
Strategy VOC Test Concept Assignment 62.0 Nearest Upsampling 61.2
Normalized Cut algorithm does not scale well with the graph size due to the generalized eigenvalue problem to solve, which hinder its use on the native image resolution (e.g., $1024\times 1024$). Thus, the clustering is applied in the latent space and yields segmentation maps at the latent resolution. To obtain pixel-level segmentation at the original image resolution, we need to upscale the low-resolution maps. In Tab. 6, we compare the nearest-neighbor upsampling approach versus our concept assignment upsampling and show that our proposed method obtain better results than the naive upsampling of the segmentation masks.
Appendix D Visual Encoders KMeans Comparison
To evaluate the potential of vision encoders for zero-shot segmentation, we compare their clustering performance with a simple KMeans algorithm. For selected vision encoders, features are extracted from the last layer and clustered with KMeans whose hyperparameter $\bm{K}$ is either determined by the ground-truth ($\bm{K}^{*}$) for each image, or fixed across the dataset. We compute the mIoU with respect to the ground truth masks using the Hungarian matching algorithm [75]. Tab. 7 shows that the diffusion encoder (SSD-1B) significantly outperforms all other vision encoders on Pascal VOC (20 classes and no background), COCO-Stuff-27 and Cityscapes. This further confirms that the diffusion features are a good candidate for localizing and segmenting objects. Interestingly, unsupervised models such as DINOv2 are better than CLIP models, suggesting that text-aligned features does not contain accurate localization features.
Model $K^{*}$ $K=3$
SSD-1B 79.5 70.8
Text-aligned
   CLIP-VIT-B/16 68.9 59.1
   CLIP-VIT-L/14 67.1 60.7
   SigLIP-B/16 62.9 55.3
   SigLIP-L/16 62.2 54.8
Unsupervised
   DINO 73.1 62.8
   DINOv2-B/14 73.8 64.8
   DINOv2-L/14 73.1 64.4
(a) VOC
Model $K^{*}$ $K=6$
SSD-1B 36.4 37.8
Text-aligned
   CLIP-VIT-B/16 31.1 31.8
   CLIP-VIT-L/14 26.4 26.6
   SigLIP-B/16 25.0 25.1
   SigLIP-L/16 22.5 23.1
Unsupervised
   DINO 33.8 34.0
   DINOv2-B/14 31.5 32.2
   DINOv2-L/14 30.9 31.5
(b) COCO-Stuff-27
Model $K^{*}$ $K=6$
SSD-1B 21.4 21.0
Text-aligned
   CLIP-VIT-B/16 14.9 14.7
   CLIP-VIT-L/14 14.3 14.0
   SigLIP-B/16 13.7 13.6
   SigLIP-L/16 11.6 11.6
Unsupervised
   DINO 18.4 17.4
   DINOv2-B/14 19.5 18.9
   DINOv2-L/14 18.2 18.1
(c) Cityscapes
Table 7: KMeans features clustering for various vision encoders.
Appendix E Hyperparameter $\tau$ tuning
We show in Sec. 4.3 that the performance of the method is highly robust with respect to the value of the threshold $\tau$. However, as the granularity of annotations varies across target datasets, the value of this threshold, fixed in our experiments, can not be in the optimal regime on each benchmark. Therefore, we relax the condition on the absence of prior knowledge about the target dataset and report in Tab. 8 results of DiffCut where $\tau$ is loosely tuned using a small subset of annotated images (200) from the target training split. Specifically, we estimate an adequate value for $\tau$ with a grid-search in the set $\{0.35,0.55,0.75\}$ for COCO-Object, COCO-Stuff and Cityscapes.
Table 8: Threshold tuning. Tuned $\tau$ is denoted with $\tau^{*}$.
DiffCut COCO-Object COCO-Stuff-27 Cityscapes $\tau=0.5$ 32.0 46.1 28.4 $\tau^{*}$ 38.7 48.6 29.8
As COCO-Object and COCO-Stuff-27 offer different level of object granularity despite corresponding to the same images, a fixed value for $\tau$ can not perform optimally on both benchmarks. Tuning the value of this threshold allows to infer the granularity of objects expected to be uncovered in images. For example, the estimated $\tau^{*}$ value for COCO-Stuff-27 is 0.35 whereas it is 0.75 on COCO-Object whose annotations requires to detect much finer objects. For Cityscapes the initial fixed $\tau$ value was in the good range to yield optimal performance.
Appendix F Image Noising
Before passing the image to the diffusion UNet, a predefined amount of gaussian noise, controlled by a parameter called the timestep, is added to it. At timestep $t=0$ the input image corresponds to the original image without added noise while $t=1000$ corresponds to an image transformed into pure gaussian noise. In Fig. 7, we show the segmentation performance on the validation split of Pascal VOC for timesteps values ranging from 0 to 500. We can observe that a small amount of noise, around 50, gives the best mIoU score, indicating that the best semantic features are obtained with a slightly noisy input image. We note that despite a significant drop in the mIoU score for $t=500$, DiffCut still reaches state-of-the-art segmentation performance on Pascal VOC benchmark, demonstrating a high robustness of the method with respect to the noising ratio.
Figure 7: mIoU according to the noising timestep on Pascal VOC.
Appendix G DiffCut Robustness
We provide in Fig. 8 an additional evaluation of the combined effect of hyperparameters $\alpha$ and $\tau$ on the segmentation performance of DiffCut for Pascal VOC dataset. Similar to Fig. 5, we observe a shift of the optimal threshold between the different curves associated to distinct $\alpha$ values. We also observe the increased robustness of DiffCut with a wide range of $\tau$ values for which the method yields mIoU scores above 60.0 for $\alpha=10$. In comparison, for $\alpha=1$, the mIoU only surpasses 60.0 between 0.91 and 0.94.
Figure 8: Robustness of DiffCut on Pascal VOC.
Appendix H Visualization of the effect of $\tau$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Ref. Image $\tau=0.3$ $\tau=0.5$ $\tau=0.7$
Figure 9: Effect of $\tau$. As $\tau$ increases, DiffCut uncover finer-grained objects.
Appendix I Semantic Coherence in Vision Encoders
Ref. Image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Ref. Image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Ref. Image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption SSD-1B DINO-VIT-B/16 DINOv2-VIT-L/14 CLIP-VIT-L/14 SigLIP-VIT-L/16
Figure 10: Qualitative results on the semantic coherence of vision encoders. We select a patch (red marker) associated to the dog in Ref. image. Top row shows the cosine similarity heatmap between the selected patch and all patches produced by vision encoders for Ref. image. Bottom row shows the heatmap between the selected patch in Ref. image and all patches produced by vision encoders for Target. image.
Appendix J Additional Visualization
Figure 11: Examples of our produced segmentation maps on COCO dataset.
Figure 12: Examples of our produced segmentation maps on Pascal Context dataset.
Appendix K Datasets Licenses
Pascal VOC:
Pascal Context:
COCO:
https://cocodataset.org/#home
Cityscapes:
https://www.cityscapes-dataset.com/
ADE20K:
https://groups.csail.mit.edu/vision/datasets/ADE20K/